### v1

@@ Abstract
Federated learning, also known as FL, is a machine learning framework in which a significant amount of clients (such as mobile devices or whole enterprises) collaborate to collaboratively train a model while keeping decentralized training data, all overseen by a central server (such as a service provider). There are advantages in terms of privacy, security, regulations, and economy with this decentralized approach to model training. FL is not impervious to the flaws that plague conventional machine learning models, despite its seeming promise. This study offers a thorough analysis of the fundamental ideas and elements of federated learning architectures, emphasizing five important areas: communication architectures, machine learning models, data partitioning, privacy methods, and system heterogeneity. We additionally address the difficulties and potential paths for future study in the area. Furthermore, based on a comprehensive review of the literature, we present a collection of architectural patterns for federated learning systems. This analysis will help to understand the basic of Federated learning, the primary components of FL, and also about several architectural details.
@@ 1 Introduction
Artificial intelligence (AI) and machine learning (ML) have gained popularity in recent years after AI’s victory over humans in the board game Alpha-Go [1]. The availability of big data and powerful processing units has accelerated the use of Machine Learning technologies in various sectors, including banking, healthcare [2], [3], [4], [5], transportation [6], customer services [7], e-commerce, and smart home applications [8]. Because machine learning techniques are so extensively employed, it is critical to ensure their security and privacy. The bulk of machine learning systems train the model by combining data from several devices or organizations on a central server or cloud platform. This is a significant disadvantage, especially when there are security threats in the training data set because to the sensitive information it includes. Several hospitals can pool their data to construct a collaborative machine-learning model for breast cancer diagnosis [9], [10] from MRI images. On the other hand, disclosing private patient data to a central server may expose confidential information to the public, which may have many negative effects. Federated Learning may be a superior choice in certain situations. Federated Learning is a cooperative learning method whereby devices or organizations exchange and aggregate the model parameters from local models rather than exchanging local data [11].
Federated Learning (FL) has profoundly impacted machine learning, particularly in terms of data security and privacy management [12]. In 2016, Google announced FL for collaborative machine learning model training across several clients, supervised by a central server [13]. Clients can range from mobile devices to entire businesses. By ensuring that the training data is decentralized, this strategy helps to reduce the hazards that come with sharing data, which is a feature of typical centralized machine learning techniques. FL has especially great promise in industries like finance and healthcare, where data protection and sensitivity are vital. The overview of federated learning is shown in Fig. 1. A federated learning system has three stakeholders: (1) the system owner, or learning coordinator; (2) the contributor client, which includes local model trainers and data contributors; and (3) the user client, which is the model user [14]. Keep in mind that a user client can also be a contributor client. System nodes, or hardware components, come in two varieties: (1) central servers; and (2) client devices.
@@
Google originally presented the concept of federated learning in 2016 when they implemented it in the Google Keyboard, enabling several Android phones to learn together. Because FL may be implemented on any edge device, it has the potential to completely transform a number of important industries, including finance, healthcare, transportation, and smart homes [15]. The most well-known instance was when scientists and doctors from various countries worked together to create an AI pandemic engine for COVID-19 diagnosis [16] using chest scans. Transportation networks present yet another intriguing use case: teaching cars to drive themselves and create city routes. In a similar vein, federated learning frameworks enable edge devices in various homes to cooperatively learn on context-aware policies for smart-home applications [17], [1].
FL functions based on two fundamental concepts: model transmission and local computing [18], [19]. Clients use their data to do local training; they only provide the trained model parameters to the central server, which combines them to update the global model [18], [20]. Until a workable model is produced, this iterative process is continued. Although FL dramatically lowers some operating expenses and systemic privacy issues, it also presents some special difficulties, such as the requirement for complex coordination and communication methods and vulnerability to a variety of attacks [21].
The authors of research paper [22] illustrate federated learning (FL) operations and highlight the distinctions between learning on a centralized data lake and on a workstation, which is shown in Fig. 2. The study [22] focuses on how federated learning (FL) may be used to incorporate data-driven machine learning (ML), particularly deep learning (DL), into medical practice. The paper examines the difficulty of exploiting massive amounts of medical data due to data silos and privacy concerns, and FL is proposed as a potential solution. Federated learning addresses privacy concerns while also enabling the creation of trustworthy and accurate machine learning (ML) models for the healthcare business without the need for centralized data collection for collaborative model training. The primary contribution of this study is to investigate federated learning (FL) as a viable solution to overcome data silos and privacy concerns in the application of ML for digital health.
@@
Figure 3 illustrates the various topologies and computation strategies that may be used to achieve a FL process. Peer-to-peer is the most preferred technique for healthcare applications, followed by an aggregate server. Because FL participants only get model parameters that are averaged among a group of participants and never have direct access to data from other institutions, FL always implicitly ensures a degree of anonymity.
@@
The focus of the research paper[23] is on Federated Learning (FL) as a distributed and privacy-preserving approach to solving wireless communication problems, especially in the context of fifth-generation (5G) networks. The study highlights the shortcomings of conventional, centralized machine learning (ML) techniques in wireless applications because of serious communication costs and worries about data privacy. Federated learning is offered as a possible solution to improve various wireless communication applications [24], [25] and avoid these problems by enabling local model training without centralizing data. The study illustrates the appropriateness of federated learning for a range of use cases by discussing numerous possible implementations of the technology within 5G networks [rahman2023icn], [26]. Federated learning, which uses locally trained models instead of gaining direct access to user data, appears to be a perfect fit for proactive caching in wireless networks, specifically for content popularity prediction, as shown in Fig. 4
@@
The aim of this study is to provide a comprehensive overview of the basic ideas and elements of federated learning architecture. We explore the key components that characterize FL systems, such as data partitioning tactics, privacy-preserving methods, the kinds of machine learning models used, communication protocols, and the management of heterogeneity within the system. Moreover, we recognize and talk about architectural patterns that provide reusable fixes for typical design issues that arise during FL system development. This review aims to improve knowledge and application of federated learning systems by summarizing ongoing research and suggesting future research avenues. In order to advance the discipline and meet the inherent issues of designing safe, reliable, and efficient federated learning systems, we hope to be of great assistance to researchers and practitioners.
@@ 2 Basic Principles of Federated Learning
Federated Learning (FL) is a machine learning approach that is based on a fundamental premise that differs from typical centralized machine learning. Among those rules are:
@@ 2.1 Localization of Data and Decentralization
2.1.1 Local Data Utilization:
FL makes sure that information is never moved to a central server, always staying on local devices or clients. Each client performs local computations and model training using its own data [27].
2.1.2 Decentralized Data Storage:
FL improves privacy and reduces the chance of data leaks by maintaining decentralized data [28]. This strategy is particularly useful in industries like healthcare and finance where data protection regulations are very strict.
2.2 Training Collaborative Models
2.2.1 Federated Training Process:
Users train together on a common global model. Each client computes model parameter updates based on its local data and sends them to a central server [29]. Apart from this, the server accommodates all the updates using raw data. Moreover, it is improving to bolster the transmission of data with minimal amount of loss. As the model iteratively moves, it adapts better to eradicate the insecure data of the various local datasets.
2.2.2 Model Aggregation:
To improve the global model, a central server compiles the updates from each client. This is an iterative process that involves multiple rounds of central aggregation and local training until the model converges. From each step, insightful insights from a wider array of data were found. This procedure allows us to get proper generalization capabilities. It is employed to train ML-based algorithms that provide precise global performance.
2.3 Maintaining Privacy
2.3.1 No Raw Data Sharing:
Federated learning significantly mitigates privacy concerns compared to centralized machine learning (ML) by enabling clients to exchange only model updates, such as gradients or parameter changes, rather than sharing raw data. This mechanism effectively reduces the risk of sensitive information being exposed during the training process. Moreover, FL empowers clients to retain control over their data, thus addressing critical issues related to data ownership and compliance with privacy regulations. By concentrating on the aggregation of model updates, federated learning can also help alleviate biases that may emerge from the consolidation of sensitive datasets. The decentralized architecture of FL not only enhances privacy but also fosters trust among participants, ultimately encouraging broader participation in collaborative training endeavors. This makes federated learning a promising approach for developing robust and privacy-preserving machine learning systems.
2.4 Strategies for Improving Privacy:
@@
2.5 Effective Communication
2.5.1 Effective contact Protocols:
To reduce latency and bandwidth utilization, FL requires frequent contact between clients and the central server [30]. As a result, effective communication protocols are crucial.
@@ 2.5.2 Model Compression and Update Sparsification:
To reduce communication costs, methods such as model compression, quantization, and update sparsification might be used.
2.6 System and Data Diversities
@@ 2.6.1 Managing Diverse Data Distributions:
The local data distributions of clients in a Federated Learning system may differ greatly if their data is non-IID (independently and identically distributed). FL algorithms need to withstand this kind of variation.
2.6.2 Client and System Variability:
The computing capacity, energy availability, and network connectivity of clients can vary. FL systems need to be built with these variances in mind.
2.7 Safety and Robustness
2.7.1 Attack Resistance:
Front-end systems (FL) need to be able to withstand a variety of attacks, including poisoning attempts, in which malevolent clients transmit destructive updates in an attempt to tamper with the global model.
2.7.2 Fault Tolerance:
Without jeopardizing the training process as a whole, the system should be able to manage client dropouts or erratic communication.
2.8 Scalable Architecture
FL systems ought to be able to effectively expand to accommodate a substantial client base, maybe reaching the millions. Scalable server-side infrastructure and effective update aggregation algorithms are needed for this.
@@ 4 Federated Learning Architectures
Federated Learning (FL) is an approach that allows machine learning models to be trained in a distributed manner using remotely hosted datasets without polluting the data through aggregation. [33]. FL is a viable way to improve ML-based systems, increase alignment with regulatory standards, and improve data sovereignty and trust. Many questions remain unanswered before FL is widely used [33]. Both federated learning and neural architecture search face many unsolved challenges. However, the search for optimal neural designs in the context of federated learning is particularly challenging [34]. This work provides background on Federated Learning and Neural Architecture Search (NAS) [34], with a particular focus on the recently developed area of Federated Neural Architecture Search (FNAS). Systems are categorized into offline and online approaches, and single- and multi-objective NAS methods are discussed. The study classifies federated learning systems, draws attention to the difficulties and limitations of online FNAS, looks at ways to balance various goals including precision and communication expenses, and concludes by summarizing the primary issues still facing FNAS.
Since its inception, federated learning has seen tremendous evolution. In 2016, Google researchers formally presented the idea, concentrating at first on enhancing user privacy for mobile keyboard prediction applications [35]. FL’s application base has grown over time to include a wide range of fields, including banking, healthcare, and other industries. The creation of novel algorithms, privacy-preserving methods [36], and designs to manage the inherent heterogeneity in federated environments are important developments.
The study [37] provides a comprehensive examination of Federated Learning (FL), emphasizing privacy-preserving solutions and focusing on enabling technologies, protocols, practical implementations, and use cases across various sectors. The paper [37] intends to help data scientists create more effective privacy-preserving solutions by offering a complete review of relevant FL protocols, platforms, and applications. It also discusses the primary advantages and disadvantages of FL, as well as detailed use cases that demonstrate how successfully it can be utilized in various industries.
@@
Vertical FL is also known as feature-based FL. Figure 8 depicts the Vertical FL procedure. In this case, data sets may differ in features but have similar sample IDs. What we are doing with Vertical FL is gathering and organizing these different elements. Next, in order to create a model that collectively incorporates data from both entities, we must compute the training loss. Every entity in Vertical FL shares the same identification and status. The Vertical FL system also presumes that its customers are trustworthy when it comes to security. Nonetheless, Vertical FL raises two security-related issues. The Vertical FL architecture consists of two primary components: encrypted model training and encrypted entity alignment [43], [44], [45].
@@
Figure 10 presents an overview of the FTL procedure. To complete the method, the Guest and Host must first compute and encrypt their findings locally. Gradients and losses are calculated using the data. They then provide Arbiter access to the encrypted values. The Arbiter then provides the Guest and Host with the gradients and loss computations, which they may use to make model adjustments. Until the loss function converges, the FTL framework iterates [46]. Additionally, FTL offers support for both homogenous and heterogeneous training methodologies. Using a variety of sample types, entities assist in training the model or models in the homogeneous method. When entities are heterogeneous, they have identical samples but differing feature spaces.
@@
Tien-Dung et al. [42] present a further FL framework. FL’s method is meant to allow for concurrent training while still protecting anonymity. A model may be trained on numerous geographically scattered training data sets—which may belong to different owners—using their framework, known as FEDF. As seen in Fig. 12, the authors’ proposed design consists of a master and X workers. Additionally, the writers were able to test their framework on a variety of systems. The major datasets utilized to assess the FEDF architecture were the CIFAR-10 membrane data set (MEMBRANE) and the health care imaging data set (HEART-VESSEL). The assessment criteria were performance, training speed, and data volume transmitted.
@@
Qiong Wu et al.’s framework for FL is another intriguing one [43]. They have focused their FL architecture on IoT adaptability. Although this study is not focused on IoT, it is worth noting that FL has been proposed for IoT. [44, 45]. Figure 13 shows the authors’ PerFit framework. PerFit was designed to help with a few FL and IoT-related issues. Upon closer inspection, PerFit’s cloud-based architecture should provide IoT devices with easily accessible processing capability, according to the authors. The architecture is set up in a way that allows any Internet of Things device to release its computational burden, thereby meeting the demands for low latency and efficiency.
@@
The creation of a privacy-preserving federated learning (FL) application for anticipating customers’ financial distress is the major focus of the study [48]. This approach addresses the resource and data privacy restrictions of traditional centralized machine learning models. The main contribution is a new FL method that allows partial task contributions, which reduces the effect of straggler agents and enhances model convergence and performance in resource-constrained contexts. The suggested approach outperforms current FL models in accuracy and maintains data privacy while achieving accuracy comparable to centralized models. The authors provided the general process flowchart in Figure 14 and provided a step-by-step breakdown of their suggested methodology as follows:

### v2

@@ Abstract
Federated Learning (FL) is a machine learning framework where multiple clients, from mobiles to enterprises, collaboratively construct a model under the orchestration of a central server but still retain the decentralized nature of the training data. This decentralized training of models offers numerous advantages, including cost savings, enhanced privacy, improved security, and compliance with legal requirements. However, for all its apparent advantages, FL is not immune to the limitations of conventional machine learning methodologies. This article provides an elaborate explanation of the inherent concepts and features found within federated learning architecture, addressing five key domains: system heterogeneity, data partitioning, machine learning models, communication protocols, and privacy techniques. This article also highlights the limitations in this domain and proposes avenues for future work. Besides, we provide a set of architectural patterns for federated learning systems, which are derived from the systematic survey of the literature. The main elements of FL, the fundamentals of Federated Learning, and a few architectural specifics will all be better understood with the aid of this research.
@@ 1 Introduction
The fields of artificial intelligence (AI) and machine learning (ML) have gained tremendous growth in the past few years, especially due to advancements in autonomous driving technologies by players like Tesla and Waymo [1]. The availability of large datasets and high-powered computing resources has enabled the use of machine learning techniques in multiple domains, such as banking, healthcare [2], [3], [4], [5], transportation [6], customer service [7], e-commerce, and smart home technologies [8]. Considering extensive use of machine learning techniques, their security and privacy is essential to protect. The majority of machine learning frameworks build the model by collecting data from different devices or organizations on a centralized server or cloud-based system. This is a significant limitation, especially in the case of security threats in the training dataset because of the sensitive nature of the data it contains. Many hospitals can combine their data to create a shared machine-learning model for detecting breast cancer [9], [10] using MRI images. On the other hand, patient data sharing with a centralized server risks the exposure of confidential data, and many negative consequences would then follow. In some situations, Federated Learning would be a preferable option. Federated Learning is a collaborative type of learning in which devices or organizations share and aggregate the model parameters of their local models, instead of local data sharing [11].
FL has profoundly impacted machine learning, particularly in terms of data security and privacy management [12]. Clients can range from mobile devices to entire businesses. By ensuring that the training data is decentralized, this strategy helps to reduce the hazards that come with sharing data, which is a feature of typical centralized machine learning techniques. FL has especially great promise in industries like finance and healthcare, where data protection and sensitivity are vital. The overview of FL is shown in Fig. 1. An FL system has three stakeholders: (1) the system owner or learning coordinator; (2) the contributor client, which includes local model trainers and data contributors; and (3) the user client, which is the model user [13]. Keep in mind that a user client can also be a contributor client. System nodes, or hardware components, come in two varieties: (1) central servers and (2) client devices.
@@
Google originally presented the concept of FL in 2016 when they implemented it in the Google Keyboard, enabling several Android phones to learn together. Because FL may be implemented on any edge device, it has the potential to completely transform a number of important industries, including finance, healthcare, transportation, and smart homes [14]. The most well-known instance was when scientists and doctors from various countries worked together to create an AI pandemic engine for COVID-19 diagnosis [15] using chest scans. Transportation networks present yet another intriguing use case: teaching cars to drive themselves and create city routes. In a similar vein, FL frameworks enable edge devices in various homes to cooperatively learn on context-aware policies for smart-home applications [16].
Therefore, when privacy, bandwidth, and regulatory compliance are important factors, FL provides a number of advantages to centralized machine learning. FL makes it possible to train models directly on local devices in industries like healthcare and finance, guaranteeing that private information stays on-site and reducing the legal issues that come with data centralization. By enabling local model training, FL lessens the need to send massive amounts of raw data in the context of the Internet of Things (IoT) and 5G applications, saving bandwidth and improving efficiency. Furthermore, by enabling businesses to create reliable models without moving personal data to central servers, FL complies with strict data protection laws like GDPR and HIPAA. Google’s Gboard is a real-world example of how FL is used; it uses FL to enhance typing predictions on users’ devices without sending private text data to central servers [17]
FL functions based on two fundamental concepts: model transmission and local computing [18]. Clients use their data to do local training; they only provide the trained model parameters to the central server, which combines them to update the global model [19], [20]. Until a workable model is produced, this iterative process is continued. Although FL dramatically lowers some operating expenses and systemic privacy issues, it also presents some special difficulties, such as the requirement for complex coordination and communication methods and vulnerability to a variety of attacks [21].
The authors of the research paper [22] illustrate (FL) operations and highlight the distinctions between learning on a centralized data lake and on a workstation, which is shown in Fig. 2. The study focuses on how FL may be used to incorporate data-driven machine learning (ML), particularly deep learning (DL), into medical practice. The paper examines the difficulty of exploiting massive amounts of medical data due to data silos and privacy concerns, and FL is proposed as a potential solution. FL addresses privacy concerns while also enabling the creation of trustworthy and accurate machine learning (ML) models for the healthcare business without the need for centralized data collection for collaborative model training. The primary contribution of this study is to investigate FL as a viable solution to overcome data silos and privacy concerns in the application of ML for digital health.
@@
Figure 3 illustrates the various topologies and computation strategies that may be used to achieve an FL process. Peer-to-peer is the most preferred technique for healthcare applications, followed by an aggregate server. Because FL participants only get model parameters that are averaged among a group of participants and never have direct access to data from other institutions, FL always implicitly ensures a degree of anonymity.
@@
The focus of the research paper[23] is on FL as a distributed and privacy-preserving approach to solving wireless communication problems, especially in the context of fifth-generation (5G) networks. The study highlights the shortcomings of conventional, centralized machine learning (ML) techniques in wireless applications because of serious communication costs and worries about data privacy. Federated learning is offered as a possible solution to improve various wireless communication applications [24], [25], and avoid these problems by enabling local model training without centralizing data. The study illustrates the appropriateness of FL for a range of use cases by discussing numerous possible implementations of the technology within 5G networks [26]. FL, which uses locally trained models instead of gaining direct access to user data, appears to be a perfect fit for proactive caching in wireless networks, specifically for content popularity prediction, as shown in Fig. 4
@@
The aim of this study is to provide a comprehensive overview of the basic ideas and elements of FL architecture. We explore the key components that characterize FL systems, such as data partitioning tactics, privacy-preserving methods, the kinds of machine learning models used, communication protocols, and the management of heterogeneity within the system. Moreover, we recognize and talk about architectural patterns that provide reusable fixes for typical design issues that arise during FL system development. This review aims to improve knowledge and application of FL systems by summarizing ongoing research and suggesting future research avenues. In order to advance the discipline and meet the inherent issues of designing safe, reliable, and efficient FL systems, we hope to be of great assistance to researchers and practitioners.
@@ 2 Basic Principles of Federated Learning
Federated Learning is a different way to use computers to learn from data. Instead of sending all their information to one place, devices like phones or computers learn on their own. They only send small updates to a main server. This keeps people’s information private and helps save money. Below are the main ideas about Federated Learning.
@@ 2.1 Localization of Data and Decentralization
2.1.1 Local Data Utilization
FL prevents sensitive data from being sent to a central server by keeping raw data on the client’s device or local system. Every client uses its own data to independently carry out local calculations and model training [27].
2.1.2 Decentralized Data Storage
FL greatly improves security and privacy by preserving decentralized data storage. This strategy lowers the possibility of data breaches and guarantees adherence to stringent laws in sectors like healthcare and finance. FL reduces these risks in contrast to centralized ML, where data aggregation raises exposure to cyberthreats [28].
2.2 Collaborative Model Training
2.2.1 Federated Training Process
FL keeps the data private while allowing for teamwork in training models on different devices or by different organizations. Each participant uses their own data to calculate and send updates to a main server. To protect your data, only encrypted updates to the model are sent, not the raw data itself. [29].
2.2.2 Model Aggregation
The central server improves a global model by collecting, processing, and combining updates from multiple clients. This process continues until the model is stable. Using techniques like Federated Averaging (FedAvg), which combines knowledge from different datasets, helps make the model more effective.
2.3 Privacy Preservation
2.3.1 No Raw Data Sharing
Federated learning helps solve privacy problems better than centralized machine learning (ML) by letting clients share only model updates, like changes to parameters or gradients, instead of raw data. This greatly lowers the chance of exposing sensitive information during training. It also allows clients to keep control of their data, addressing important issues related to data ownership and privacy laws. By focusing on combining model updates, federated learning can reduce biases that may come from merging sensitive datasets. The decentralized structure of federated learning not only boosts privacy but also builds trust among participants, encouraging more people to take part in collaborative training. This makes federated learning a useful way to create strong and privacy-friendly machine learning systems.
2.4 Strategies for Improving Privacy
@@
2.5 Communication Efficiency
2.5.1 Optimized Communication Protocols
In FL, efficient communication protocols are crucial for reducing latency and bandwidth usage. Protocols that are optimized minimize synchronization delays and provide seamless communication between clients and the central server [30].
@@ 2.5.2 Model Compression and Update Sparsification:
Due to the possibility of high communication costs from frequent model updates, FL uses compression techniques like:
  • Quantization: Reduces the precision of model parameters to lower data transmission costs
  • Update sparsification: Reduces the amount of data that is transmitted by choosing only the most important parameter changes.
2.6 Handling Data and System Heterogeneity
@@ 2.6.1 Managing Diverse Data Distributions:
Client data in federated learning applications often varies significantly and is not uniform. This can harm the performance of the overall model because different clients may have very different data patterns. FL algorithms must be adjusted to handle these variations effectively.
2.6.2 Addressing Client and System Variability
Many client devices have different processing abilities, energy levels, and network conditions that must be supported by federated learning systems. Techniques like asynchronous aggregation and customized training schedules help make it easier for devices with limited resources to participate effectively.
2.7 Security and Robustness
2.7.1 Defense Against Model Poisoning Attacks
Adversarial threats such as model poisoning, in which malevolent clients introduce modified updates to taint the global model, must be thwarted by FL. Such risks are lessened by methods like anomaly detection and Byzantine-resilient aggregation.
2.7.2 Fault Tolerance and Resilience
Unreliable network connections and client dropouts must be supported by FL systems. Techniques like redundant calculations and backup models guarantee that the system is robust even if certain clients do not show up for training sessions.
2.8 Scalability Considerations
FL needs to be built with efficiency in mind, supporting millions of clients. Important scalability enhancements consist of:
  • Hierarchical FL: It distributes the workload by introducing intermediary aggregation nodes between the clients and the central server.
  • Edge computing integration: Permits edge servers to submit updates to the cloud after completing partial aggregation.
3 Explainable AI (XAI) and Zero Trust Architecture (ZTA)
Integrating Explainable AI (XAI) into Zero Trust Architecture (ZTA) is essential for maintaining security, accountability, and trust in federated learning (FL) systems. ZTA follows the principle of ”never trust, always verify,” which requires continuous re-evaluation of authentication and access decisions. XAI provides transparency by explaining why requests are approved or denied and identifying the causes of unusual behavior, ensuring compliance with security policies.
For example, if a hospital node is flagged for suspicious data, XAI can reveal specific features, like abnormal timestamps or value ranges, helping auditors understand the alert. In finance, XAI enhances transaction monitoring by highlighting unusual transfer amounts or geolocation patterns, helping to detect fraud. [31].
3.1 XAI for Access Control and Anomaly Detection
In a zero-trust, federated environment, interpretability is crucial. When a user or device requests access, an explainable AI (XAI) system clarifies the decision-making process. It can reveal whether access was denied due to the device’s security posture or if the user’s authentication history warranted extra scrutiny. By presenting this information in clear terms, XAI turns complex decisions into actionable insights.
For anomaly detection, XAI models analyze large volumes of network data to identify deviations from established baselines, providing contextual explanations for these anomalies. This allows security analysts to quickly understand issues, whether it’s unusual traffic to an unknown port or unexpected device communication.
Additionally, these explanations are logged, creating a verifiable record that proves compliance with Zero Trust Architecture (ZTA) principles. In 5G network security, for example, combining federated learning (FL) with XAI enables operators to learn normal traffic behaviors in real time and quickly identify and explain any deviations, catching intrusions early. [32].
3.2 Federated Learning’s Role in Zero Trust Architecture (ZTA)
Federated Learning supports Zero Trust principles by decentralizing model training and reducing the reliance on a central data store. Instead of collecting sensitive information in one place, it lets each node like a hospital database or edge device train locally and only send updates about the model. This method improves data confidentiality, lowers the risk of attacks, and establishes strict controls over sensitive information.
Additionally, real-time behavior analysis helps adjust trust levels in the system. Nodes that follow security rules earn high trust scores, while those that act suspiciously lose trust. Overall, the ongoing learning process of Federated Learning, together with Explainable Artificial Intelligence (XAI), aligns with the focus of Zero Trust Architecture (ZTA) on continuous verification and limited access [33].
@@ 5 Federated Learning Architectures
FL is an approach that allows machine learning models to be trained in a distributed manner using remotely hosted datasets without polluting the data through aggregation. [36]. FL is a viable way to improve ML-based systems, increase alignment with regulatory standards, and improve data sovereignty and trust. Several challenges remain, including data heterogeneity, privacy risks, communication efficiency, and secure model aggregation, which must be addressed before FL achieves widespread adoption. Both federated learning and neural architecture search face many unsolved challenges. However, the search for optimal neural designs in the context of federated learning is particularly challenging [37]. This work provides background on Federated Learning and Neural Architecture Search (NAS), with a particular focus on the recently developed area of Federated Neural Architecture Search (FNAS). Systems are categorized into offline and online approaches, and single- and multi-objective NAS methods are discussed. The study classifies federated learning systems, draws attention to the difficulties and limitations of online FNAS, looks at ways to balance various goals including precision and communication expenses, and concludes by summarizing the primary issues still facing FNAS.
FL’s application base has grown over time to include a wide range of fields, including banking, healthcare, and other industries. The creation of novel algorithms, privacy-preserving methods [38], and designs to manage the inherent heterogeneity in federated environments are important developments.
The study [39] provides a comprehensive examination of FL, emphasizing privacy-preserving solutions and focusing on enabling technologies, protocols, practical implementations, and use cases across various sectors. It intends to help data scientists create more effective privacy-preserving solutions by offering a complete review of relevant FL protocols, platforms, and applications. It also discusses the primary advantages and disadvantages of FL, as well as detailed use cases that demonstrate how successfully it can be utilized in various industries.
@@
Vertical FL is also known as feature-based FL. Figure 8 depicts the Vertical FL procedure. In this case, data sets may differ in features but have similar sample IDs. What we are doing with Vertical FL is gathering and organizing these different elements. Next, in order to create a model that collectively incorporates data from both entities, we must compute the training loss. Every entity in Vertical FL shares the same identification and status. The Vertical FL system also presumes that its customers are trustworthy when it comes to security. Nonetheless, Vertical FL raises two security-related issues. The Vertical FL architecture consists of two primary components: encrypted model training and encrypted entity alignment [46], [47].
@@
Figure 10 presents an overview of the FTL procedure. To complete the method, the Guest and Host must first compute and encrypt their findings locally. Gradients and losses are calculated using the data. They then provide Arbiter access to the encrypted values. The Arbiter then provides the Guest and Host with the gradients and loss computations, which they may use to make model adjustments. Until the loss function converges, the FTL framework iterates. Additionally, FTL offers support for both homogenous and heterogeneous training methodologies. Using a variety of sample types, entities assist in training the model or models in the homogeneous method. When entities are heterogeneous, they have identical samples but differing feature spaces.
@@
FL’s method is meant to allow for concurrent training while still protecting anonymity. A model may be trained on numerous geographically scattered training data sets—which may belong to different owners—using their framework, known as FEDF. As seen in Fig. 12, the authors’ proposed design consists of a master and X workers. Additionally, the writers were able to test their framework on a variety of systems. The major datasets utilized to assess the FEDF architecture were the CIFAR-10 membrane data set (MEMBRANE) and the health care imaging data set (HEART-VESSEL). The assessment criteria were performance, training speed, and data volume transmitted.
@@
Figure 13 shows the authors’ PerFit framework. PerFit was designed to help with a few FL and IoT-related issues. Upon closer inspection, PerFit’s cloud-based architecture should provide IoT devices with easily accessible processing capability, according to the authors. The architecture is set up in a way that allows any Internet of Things device to release its computational burden, thereby meeting the demands for low latency and efficiency.
@@
The creation of a privacy-preserving FL application for anticipating customers’ financial distress is the major focus of the study [50]. This approach addresses the resource and data privacy restrictions of traditional centralized machine learning models. The main contribution is a new FL method that allows partial task contributions, which reduces the effect of straggler agents and enhances model convergence and performance in resource-constrained contexts. The suggested approach outperforms current FL models in accuracy and maintains data privacy while achieving accuracy comparable to centralized models. The authors provided the general process flowchart in Figure 14 and provided a step-by-step breakdown of their suggested methodology as follows: